# Subspace State Transfer in Coined Quantum Walks

### v1

@@ 9 Infinite families
In this section, we construct infinite families of Grover coined walks that admit perfect pointwise $d$-dimensional subspace state transfer between the marked vertices, where $d\geq 2$. We will assume $C_{a}=C_{b}$. Let $W$ be a subspace of $\operatorname{col}(C_{a}+I)$. We consider two special cases that make the analysis easier.
@@
Note that 0 is always an eigenvalue of $G$, and its eigenspace contains
$e_{1}-e_{\deg(a)+1},\;e_{2}-e_{\deg(a)+2},\;\cdots,\;e_{\deg(a)}-e_{2\deg(a)}.$
Thus if $x_{a}(W)$ is $\gamma$-strongly cospectral to $x_{b}(W)$, then they are strongly cospectral. Moreover, any eigenvector $x$ for $G$ orthogonal to these vectors satisfies $x[\operatorname{cl}(a)]=x[\operatorname{cl}(b)]$. It follows that strong cospectrality between $x_{a}(W)$ and $x_{b}(W)$ depends only on the projections of the column spaces of $E_{0}[\operatorname{cl}(a),\operatorname{cl}(a)]$ and $E_{0}[\operatorname{cl}(b),\operatorname{cl}(b)]$ onto $W$. For ease of notation, partition $A(X\setminus\{a,b\})$ as
@@
where $A_{1}$ is indexed by $N(a)$, and $A_{3}$ is indexed by the remaining vertices in $X\setminus\{a,b\}$.
@@
Let $U=RC$ be the transition matrix of a Grover coined walk on $X$ with marked vertices $a$ and $b$. Suppose $C_{a}=C_{b}$. Let $W$ be a subspace of $\operatorname{col}(C_{a}+I)$. Let
@@
be the spectral decomposition of the $(C_{a},C_{b})$-expansion of $X$. There is perfect pointwise $W$-transfer from $a$ to $b$ if and only if $W$ is orthogonal to the column space of
@@
    $0=\cos(j\pi/t)$ for some even $j$, and any $\lambda$ for which the column space of $E_{\lambda}[\operatorname{cl}(b),\operatorname{cl}(b)]$ is not orthogonal to $W$ can be written as
    $\lambda=\cos(j\pi/t)$
    for some odd $j$.
@@
    $0=\cos(j\pi/t)$ for some odd $j$, and any $\lambda$ for which the column space of $E_{\lambda}[\operatorname{cl}(b),\operatorname{cl}(b)]$ is not orthogonal to $W$ can be written as
    $\lambda=\cos(j\pi/t)$
    for some even $j$.
Proof. It follows from Theorem 8.2 that the $x_{a}(W)$ and $x_{b}(W)$ are strongly cospectral if and only if the projections of $E_{0}[\operatorname{cl}(a),\operatorname{cl}(a)]$ and $E_{0}[\operatorname{cl}(b),\operatorname{cl}(b)]$ onto $W$ are either equal or opposite. By Lemma 8.3, this requires the projection of every column of
$\begin{pmatrix}A_{1}&A_{2}\end{pmatrix}\Delta^{-1/2}E_{0}[V(X)\setminus\{a,b\},V(X)\setminus\{a,b\}]$
onto $W$ to be zero. Moreover, if
$\operatorname{proj}_{W}\!\left(E_{0}[\operatorname{cl}(a),\operatorname{cl}(a)]\right)=\operatorname{proj}_{W}\!\left(E_{0}[\operatorname{cl}(b),\operatorname{cl}(b)]\right),$
then 0 is the only eigenvalue lying in $\Lambda_{a(W),b(W)}^{+}$, and if
$\operatorname{proj}_{W}\!\left(E_{0}[\operatorname{cl}(a),\operatorname{cl}(a)]\right)=-\operatorname{proj}_{W}\!\left(E_{0}[\operatorname{cl}(b),\operatorname{cl}(b)]\right),$
then 0 is the only eigenvalue lying $\Lambda_{a(W),b(W)}^{-}$.
We apply this characterization to construct two infinite families of walks that admit perfect pointwise subspace state transfer. The first family is able to transfer any subspace fixed by any reflection coin assigned to the marked vertices on $K_{2,m}$.
@@
Let $a$ and $b$ be the two vertices in $K_{2,m}$ of degree $m$. Let $U=RC$ be the transition matrix of a Grover coined walk on $K_{2,m}$ with marked vertices $a$ and $b$ and any reflection coins $C_{a}=C_{b}$. For any subspace $W$ of $\operatorname{col}(C_{a}+I)$, there is perfect pointwise $W$-transfer from $a$ to $b$ at time $t=2$.
Proof. Since $A\setminus\{a,b\}=0$, its column space is orthogonal to $W$ for any subspace $W$ of $\operatorname{col}(C_{a}+I)$. Moreover,
$G\begin{pmatrix}x\\ x\\ z\end{pmatrix}=\lambda\begin{pmatrix}x\\ x\\ z\end{pmatrix}$
if and only if
$\begin{pmatrix}0&P\\ P&0\end{pmatrix}\begin{pmatrix}x\\ z\end{pmatrix}=\lambda\begin{pmatrix}x\\ z\end{pmatrix}$,
where $P$ is the projection onto $\operatorname{col}(C_{a}+I)$. Since the non-zero eigenvalue of
$\begin{pmatrix}0&P\\ P&0\end{pmatrix}$
are $\pm 1$ with eigenprojections
$\frac{1}{2}\begin{pmatrix}P&P\\ P&P\end{pmatrix},\quad\frac{1}{2}\begin{pmatrix}P&-P\\ -P&P\end{pmatrix}$
we have
@@
It follows from Theorem 9.1 that perfect pointwise $W$-transfer occurs from $a$ to $b$ at time $2$.
@@
For any positive integer $m$, let $c$ and $d$ be two distinct integers between 0 and $m$ such that $c+d=m$. Let $U=RC$ be the transition matrix of a Grover coined walk on $X(2m,\pm\{c,d\})$ with marked vertices 0 and $m$ and
$C_{0}=C_{m}=I-\frac{1}{2}J.$
Let
$W=\mathrm{span}\left\{\begin{pmatrix}1\\ 0\\ -1\\ 0\end{pmatrix},\begin{pmatrix}0\\ 1\\ 0\\ -1\end{pmatrix}\right\}$.
There is perfect pointwise $W$-transfer from 0 to $m$ at time $t=4$.
Proof. Let $a=0$ and $b=m$. Then $N(a)=N(b)$. Since c+c=ddc+c=-d-d and cd=d+$c$-d=-d+c in ${\mathbb{Z}}_{2m}$, the vertices $c$ and d-d are twins in $X\setminus\{a,b\}$, and so are the vertices c-c and $d$. Thus $A_{1}=0$, and $W$ is orthogonal to $\operatorname{col}(A_{2})$.
It suffices to check the parity condition for the eigenvalues. A calculation shows that for any vector $w\in W$,
$G\begin{pmatrix}w\\ w\\ 0\end{pmatrix}=\frac{1}{\sqrt{2}}\begin{pmatrix}w\\ w\\ 0\end{pmatrix},\quad G\begin{pmatrix}w\\ -w\\ 0\end{pmatrix}=-\frac{1}{\sqrt{2}}\begin{pmatrix}w\\ -w\\ 0\end{pmatrix}$.
Thus any eigenvector for $G$ orthogonal to these eigenvectors must have its $N(a)$ and $N(b)$ components orthogonal to $W$. Therefore,
$\Lambda^{+}_{a(W),b(W)}=\{0\}=\{\cos(2\pi/4)\}$
@@
$\Lambda^{-}_{a(W),b(W)}=\{\pm 1/\sqrt{2}\}=\{\cos(\pi/4),\quad\cos(3\pi/4)\}.$
It follows from Theorem 9.1 that perfect pointwise $W$-transfer occurs from $a$ to $b$ at time $4$.
@@
Let $X=GP(k,n)$ with $a$ and $b$ being the vertices of degree $k$. Let $U=RC$ be the transition matrix of a Grover coined walk on $X$ with marked vertices $a$ and $b$ and any reflection coin $C_{a}=C_{b}$. Let $W$ be any subspace of $\operatorname{col}(C_{a}+I)$. There is perfect pointwise $W$-transfer from $a$ to $b$ at time n1n-1.
@@
[…]
Note that for any vector $w\in\operatorname{col}(P)$, the vector $z\otimes w$ lies in the kernel of $\lambda^{2}I+\lambda B+FF^{*}$ if and only if the entries $z_{j}$ of $z$ satisfy the reccurence relation
$z_{j+1}=2\lambda z_{j}-z_{j-1},\quad j=2,3,\cdots,m-1$
with boundary contitions
$(1-2\lambda^{2})x_{1}+\lambda z_{2}=0$
$\lambda x_{m-1}+(1-2\lambda^{2})z_{m}=0.$
Solving this recurrence relation yields
$\lambda=\cos\frac{\ell\pi}{m+1},\qquad z=\begin{pmatrix}\cos\frac{\ell\pi}{m+1}\\ \vdots\\ \cos\frac{j\ell\pi}{m+1}\\ \vdots\\ \cos\frac{m\ell\pi}{m+1}\end{pmatrix},\qquad\ell=0,1,\cdots,m$ (13)
On the other hand, for any vector $w\in\ker(P)$, the vector $z\otimes w$ lies in the kernel of $\lambda^{2}I+\lambda B+FF^{*}$ if and only if
$\lambda I\otimes(-2\lambda I+A(P_{m}))z=0.$
Using eigenvectors of $A(P_{m})$ we obtain solutions
$\lambda=\cos\frac{\ell\pi}{m+1},\qquad z=\begin{pmatrix}\sin\frac{\ell\pi}{m+1}\\ \vdots\\ \sin\frac{j\ell\pi}{m+1}\\ \vdots\\ \sin\frac{m\ell\pi}{m+1}\end{pmatrix},\qquad\ell=0,1,\cdots,m$ (14)
By Lemma 8.3, Equations (13) and (14) provide $k(m+1)$ orthogonal eigenvectors for $G$ with corresponding $\lambda$ as eigenvalues. Extend these to a basis of orthogonal eigenvectors for $G$. It follows that only those eigenvectors $x$ arising from Equation (13) satisfy
$\operatorname{proj}_{W}\!\left(x[\operatorname{cl}(a)]\right)=\operatorname{proj}_{W}\!\left(x[\operatorname{cl}(a)]\right)\neq 0.$
Moreover, since
$\operatorname{col}\frac{m\pi}{m+1}=\cos\left(\ell\pi-\frac{\ell\pi}{m+1}\right),$
whenever $\ell$ is even we have
$\operatorname{proj}_{W}\!\left(x[\operatorname{cl}(a)]\right)=\operatorname{proj}_{W}\!\left(x[\operatorname{cl}(b)]\right),$
and whenever $\ell$ is odd we have
$\operatorname{proj}_{W}\!\left(x[\operatorname{cl}(a)]\right)=-\operatorname{proj}_{W}\!\left(x[\operatorname{cl}(a)]\right).$
Thus by Theorem 9.1, perfect pointwise $W$-transfer occurs from $a$ to $b$ at time $m+1$.

### v2

@@ 9 Infinite families
In this section, we construct infinite families of Grover coined walks that admit pointwise perfect or pretty good $d$-dimensional subspace state transfer between the marked vertices, where $d\geq 2$. We will assume $C_{a}=C_{b}$. Let $W$ be a subspace of $\operatorname{col}(C_{a}+I)$. We consider two special cases that make the analysis easier.
@@
Note that if $C_{a}=C_{b}$, then 0 is always an eigenvalue of $G$, and its eigenspace contains
$e_{1}-e_{\deg(a)+1},\;e_{2}-e_{\deg(a)+2},\;\cdots,\;e_{\deg(a)}-e_{2\deg(a)}.$ (17)
Therefore, any eigenvector $z$ for $G$ orthogonal to these vectors must satisfy $z[\operatorname{cl}(a)]=z[\operatorname{cl}(b)]$. It follows that $x_{a}(W)$ is $\gamma$-strongly-cospectral to $x_{b}(W)$ if and only if it is strongly-cospectral to $x_{b}(W)$, which holds precisely when every vector $z$ in $\ker(G)$ satisfies
$\operatorname{proj}_{W}\!\left(z[\operatorname{cl}(a)]\right)=-\operatorname{proj}_{W}\!\left(z[\operatorname{cl}(b)]\right).$
For ease of notation, partition $A(X\setminus\{a,b\})$ as
@@
where $A_{1}$ is indexed by $N(a)$, and $A_{3}$ is indexed by the other vertices in $X\setminus\{a,b\}$.
@@
Let $U=RC$ be the transition matrix of a Grover coined walk on $X$ with marked vertices $a$ and $b$. Suppose $a$ and $b$ are twins and $C_{a}=C_{b}$. Let $W$ be a subspace of $\operatorname{col}(C_{a}+I)$. Let
@@
be the spectral decomposition of the $(C_{a},C_{b})$-expansion of $X$. Pointwise perfect $W$-transfer occurs from $a$ to $b$ at time $t$ if and only if $W$ is orthogonal to the column space of
@@
    $t$ is a multiple of $4$, and $t\arccos\lambda$ is an odd multiple of $\pi$ for each $\lambda\in\Lambda_{a(W)}\setminus\{0\}$.
@@
    $t\equiv 2\pmod{4}$, and $t\arccos\lambda$ is an even multiple of $\pi$ for each $\lambda\in\Lambda_{a(W)}\setminus\{0\}$.
Proof. Let $z$ be any 0-eigenvector for $G$. Let $x$ and $y$ be the following components of $z$:
$x=z[\operatorname{cl}(a)\cup\operatorname{cl}(b)],\quad y=z[V(X)\setminus\{a,b\}].$
By Lemma 8.4,
$\begin{pmatrix}\frac{C_{a}+I}{2}&\frac{C_{b}+I}{2}\\ 0&0\end{pmatrix}x=-A(X\setminus\{a,b\})\Delta^{-\frac{1}{2}}y.$
Let $P_{W}$ be the orthogonal projection onto $W$. Since $C_{a}=C_{b}$ and $W$ is a subspace of $\operatorname{col}(C_{a}+I)$,
$P_{W}{x[\operatorname{cl}(a)]}+P_{W}{x[\operatorname{cl}(b)]}$ $=\begin{pmatrix}P_{W}&0\end{pmatrix}\begin{pmatrix}\frac{C_{a}+I}{2}&\frac{C_{b}+I}{2}\\ 0&0\end{pmatrix}x$
$=-\begin{pmatrix}P_{W}&0\end{pmatrix}A(X\setminus\{a,b\})\Delta^{-\frac{1}{2}}y$
$=-P_{W}{\begin{pmatrix}A_{1}&A_{2}\end{pmatrix}\Delta^{-\frac{1}{2}}y}.$
It follows from our earlier discussion that $x_{a}(W)$ and $x_{b}(W)$ are strongly-cospectral if and only if the 0-eigenprojection $E_{0}$ of $G$ satisfies
$P_{W}\begin{pmatrix}A_{1}&A_{2}\end{pmatrix}\Delta^{-\frac{1}{2}}E_{0}[V(X)\setminus\{a,b\},V(X)\setminus\{a,b\}]=0.$
In this case, 0 is the only eigenvalue in $\Lambda_{a(W),b(W)}^{-}$, and since $\arccos 0=\pi/2$, pointwise perfect $W$-transfer occurs from $a$ to $b$ if and only if one of the parity conditions in (i) and (ii) holds.
We define the support of a subspace $W$ to be
$\operatorname{supp}(W)=\bigcup_{w\in W}\operatorname{supp}(w).$
The following result describes $\Lambda_{a(W),b(W)}^{+}$ and $\Lambda_{a(W),b(W)}^{-}$ for strongly-cospectral $x_{a}(W)$ and $x_{b}(W)$ when all vertices in $\operatorname{supp}(W)$ have the same degree.
9.2 Corollary.
Let $U=RC$ be the transition matrix of a Grover coined walk on $X$ with marked vertices $a$ and $b$. Suppose $a$ and $b$ are twins and $C_{a}=C_{b}$. Let
$W\subseteq\operatorname{col}(C_{a}+I)\cap\ker\begin{pmatrix}A_{1}\\ A_{2}^{T}\end{pmatrix}$.
Then $x_{a}(W)$ and $x_{b}(W)$ are strongly-cospectral. Moreover, if $W$ is supported by vertices of the same degree $\delta$, then
$\Lambda_{a(W),b(W)}^{+}=\left\{\pm\sqrt{\frac{2}{\delta}}\right\},\quad\Lambda_{a(W),b(W)}^{-}=\{0\}.$
Proof. Since $W$ is orthogonal to the column space of $\begin{pmatrix}A_{1}&A_{2}\end{pmatrix}$, by Theorem 9.1, $x_{a}(W)$ and $x_{b}(W)$ are strongly-cospectral, and $\Lambda_{a(W),b(W)}^{-}=\{0\}$.
Let $P$ and $P_{W}$ be the projections onto $\operatorname{col}(C_{a}+I)$ and $W$, respectively. Let $\lambda$ be any nonzero eigenvalue of $G$. From Lemma 8.4 we see that
$G\begin{pmatrix}x\\ y\end{pmatrix}=\lambda\begin{pmatrix}x\\ y\end{pmatrix}$
if and only if
$\begin{pmatrix}P&0\\ P&0\end{pmatrix}\Delta^{-1/2}y=\lambda x$ (18)
and
$\left(\lambda^{2}I-\lambda\Delta^{-1/2}A(X\setminus\{a,b\})\Delta^{-1/2}-\Delta^{-1/2}\begin{pmatrix}2P&0\\ 0&0\end{pmatrix}\Delta^{-1/2}\right)y=0.$ (19)
Pick any nonzero $w\in W$. Since all vertices in $\operatorname{supp}(W)$ have degree $\delta$,
$\Delta^{-1/2}\begin{pmatrix}w\\ 0\end{pmatrix}=\delta^{-1/2}\begin{pmatrix}w\\ 0\end{pmatrix}$.
Set $y=\begin{pmatrix}w\\ 0\end{pmatrix}$ and plug it into the left-hand side of Equation (19). We get
$\left(\lambda^{2}I-\lambda\Delta^{-1/2}A(X\setminus\{a,b\})\Delta^{-1/2}-\Delta^{-1/2}\begin{pmatrix}2P&0\\ 0&0\end{pmatrix}\Delta^{-1/2}\right)y$
$=$ $\lambda^{2}\begin{pmatrix}w\\ 0\end{pmatrix}-\lambda\delta^{-1/2}\Delta^{-1/2}\begin{pmatrix}A_{1}\\ A_{2}^{T}\end{pmatrix}w-2\delta^{-1}\begin{pmatrix}w\\ 0\end{pmatrix}$
$=$ $(\lambda^{2}-2\delta^{-1})\begin{pmatrix}w\\ 0\end{pmatrix}$,
which is zero if and only if $\lambda=\pm\sqrt{2}/\sqrt{\delta}$. By Equation (18),
$\begin{pmatrix}\pm w/\sqrt{2}\\ \pm w/\sqrt{2}\\ w\\ 0\end{pmatrix}$
is an eigenvector for $G$ with eigenvalue $\lambda=\pm\sqrt{2}/\sqrt{\delta}$. Thus for any $\mu$-eigenvector $z$ with $\mu\neq\pm\sqrt{2}/\sqrt{\delta}$, the component $z[N(a)]$ must be orthogonal to $W$. It follows from Equation (18) that $\mu$ does not lie in $\Lambda_{a(W),b(W)}^{+}$. Therefore, $\Lambda_{a(W),b(W)}^{+}=\{\pm\sqrt{2}/\sqrt{\delta}\}$.
We use this observation to construct three infinite families of walks that admit pointwise perfect subspace state transfer. The first family is able to transfer any subspace fixed by any reflection coin assigned to the marked vertices on $K_{2,m}$.
@@
Let $a$ and $b$ be the two vertices in $K_{2,m}$ of degree $m$. Let $U=RC$ be the transition matrix of a Grover coined walk on $K_{2,m}$ with marked vertices $a$ and $b$ and any reflection coins $C_{a}=C_{b}$. For any subspace $W$ of $\operatorname{col}(C_{a}+I)$, there is pointwise perfect $W$-transfer from $a$ to $b$ at time $t=2$.
Proof. Note that $A\setminus\{a,b\}=0$ and the zero subspace is orthogonal to any subspace $W$ of $\operatorname{col}(C_{a}+I)$. Since all neighbors of $a$ have degree $2$, by Corollary 9.2, $x_{a}(W)$ and $x_{b}(W)$ are strongly-cospectral with
@@
It follows from Theorem 9.1 (ii) that pointwise perfect $W$-transfer occurs from $a$ to $b$ at time $2$.
@@
For any positive integer $m$, let $c$ and $d$ be two distinct integers between 0 and $m$ such that $c+d=m$. Let $U=RC$ be the transition matrix of a Grover coined walk on $X(2m,\pm\{c,d\})$ with marked vertices 0 and $m$. Let $W$ be the subspace of ${\mathbb{C}}^{\{c,d,-d,-c\}}$ spanned by
$\left\{\begin{pmatrix}1\\ 0\\ -1\\ 0\end{pmatrix},\begin{pmatrix}0\\ 1\\ 0\\ -1\end{pmatrix}\right\}$.
Let $C_{0}=C_{m}$ be the reflection about any subspace containing $W$. There is pointwise perfect $W$-transfer from 0 to $m$ at time $t=4$.
Proof. Let $a=0$ and $b=m$. First note that $a$ and $b$ are twins in $X$ as $c+d=m$:
$N(a)=\{c,d,-d,-c\}=N(b).$
Next, since 2c=2d2c=-2d in ${\mathbb{Z}}_{2m}$, the vertices $c$ and d-d are twins in $X\setminus\{a,b\}$:
$N(c)=\{2c,m,c-d,0\}=N(-d)$
and so are the vertices c-c and $d$:
$N(d)=\{m,2d,0,d-c\}=N(-c).$
Therefore $W$ is orthogonal to $\operatorname{col}\begin{pmatrix}A_{1}&A_{2}\end{pmatrix}$. Finally, since $X$ is $4$-regular, by Corollary 9.2, $x_{a}(W)$ and $x_{b}(W)$ are strongly-cospectral with
$\Lambda^{+}_{a(W),b(W)}=\{\pm 1/\sqrt{2}\}=\{\cos(\pi/4),\cos(3\pi/4)\},$
@@
$\Lambda^{-}_{a(W),b(W)}=\{0\}=\{\cos(2\pi/4)\}.$
It follows from Theorem 9.1 (i) that pointwise perfect $W$-transfer occurs from $a$ to $b$ at time $4$.
A double cone over $X$, denoted ${\mkern 2.0mu\overline{\mkern-2.0muK_{2}}}+X$, is a graph obtained from $X$ by adding two vertices, called the conical vertices, and joining them to all vertices in $X$. The following result shows a double cone over the disjoint union of $k$ cycles, whose lengths are all divisible by $4$, can transfer a $2k$-dimensional subspace between the two marked conical vertices.
9.5 Theorem.
Let $X$ be the double cone
$X={\mkern 2.0mu\overline{\mkern-2.0muK_{2}}}+(C_{4m_{1}}\cup C_{4m_{2}}\cup\cdots\cup C_{4m_{k}})$
where $m_{1},m_{2},\ldots,m_{k}$ are some positive integers. Let $a$ and $b$ be the conical vertices. Let $W$ be the subspace of functions on $V(X)\setminus\{a,b\}$ spanned by those that take alternating values 1,0,1,01,0,-1,0 along some cycles $C_{4m_{j}}$ and vanish on all other cycles. Let $C_{a}=C_{b}$ be the reflection about any subspace containing $W$. There is pointwise perfect $W$-transfer from $a$ to $b$ at time $t=4$.
Proof. Any function that takes alternating values 1,0,1,01,0,-1,0 along $C_{4m_{j}}$ lies in the kernel of $A(C_{4m_{j}})$. Hence, $W$ is orthogonal to the column space of $A(X\setminus\{a,b\})$. Since all neighbors of $a$ have degree $4$, by Corollary 9.2, $x_{a}(W)$ and $x_{b}(W)$ are strongly-cospectral with
$\Lambda^{+}_{a(W),b(W)}=\{\pm 1/\sqrt{2}\}=\{\cos(\pi/4),\cos(3\pi/4)\},$
and
$\Lambda^{-}_{a(W),b(W)}=\{0\}=\{\cos(2\pi/4)\}.$
It follows from Theorem 9.1 (i) that pointwise perfect $W$-transfer occurs from $a$ to $b$ at time $4$.
If we relax to pointwise pretty good subspace state transfer, we obtain more examples from double cones. To start, we cite a useful result about geodetic angles. A real number is a pure geodetic angle if any one of its six squared trigonometric functions is either rational or infinite.
9.6 Theorem.
[2, Ch 11] If $q\in{\mathbb{Q}}$ and $q\pi$ is a pure geodetic angle, then
$\tan(q\pi)\in\left\{0,\pm{\sqrt{3}},\pm\frac{1}{\sqrt{3}},\pm 1\right\}.$
9.7 Theorem.
Let $Z$ be a $k$-regular graph with singular adjacency matrix. Let $W$ be a subspace of $\ker(A(Z))$. Let $X={\mkern 2.0mu\overline{\mkern-2.0muK_{2}}}+Z$ with conical vertices $a$ and $b$. Let $C_{a}=C_{b}$ be the reflection about any subspace containing $W$. If $k\notin\{0,2,6\}$, then pointwise pretty good $W$-transfer occurs from $a$ to $b$ in integer steps.
Proof. By Corollary 9.2 we see that $x_{a}(W)$ and $x_{b}(W)$ are strongly-cospectral, with
$\Lambda_{a(W),b(W)}^{+}=\left\{\pm\sqrt{\frac{2}{k+2}}\right\},\quad\Lambda_{a(W),b(W)}^{-}=\{0\}.$
It suffices to show Theorem 8.2 (ii) holds for $\gamma=-1$. Let $\ell_{0}$, $\ell_{+}$ and $\ell_{-}$ be integers such that
$\ell_{0}\arccos 0+\ell_{+}\arccos\sqrt{\frac{2}{k+2}}+\ell_{-}\arccos\left(-\sqrt{\frac{2}{k+2}}\right)\equiv 0\pmod{2\pi}$.
Then
$\left(\frac{\ell_{0}}{2}+\ell_{-}\right)\pi+(\ell_{+}-\ell_{-})\arccos\sqrt{\frac{2}{k+2}}\equiv 0\pmod{2\pi}$.
If $\ell_{+}\neq\ell_{-}$, then
$\arccos\sqrt{\frac{2}{k+2}}=q\pi$
for some $q\in{\mathbb{Q}}$, and $q\pi$ would be pure geodetic. However, Theorem 9.6 implies this is impossible for $k\notin\{0,2,6\}$. Therefore $\ell_{+}=\ell_{-}$, and we have
$\sum_{\lambda\in\Lambda_{a(W),b(W)}^{+}}\ell_{\lambda}=\ell_{+}+\ell_{-}\equiv 0\pmod{2}$.
@@
Let $X=GP(k,n)$ with $a$ and $b$ being the vertices of degree $k$. Let $U=RC$ be the transition matrix of a Grover coined walk on $X$ with marked vertices $a$ and $b$ and any reflection coin $C_{a}=C_{b}$. Let $W$ be any subspace of $\operatorname{col}(C_{a}+I)$. There is pointwise perfect $W$-transfer from $a$ to $b$ at time n1n-1.
@@
$G=\begin{pmatrix}0&\frac{1}{\sqrt{2}}Q&0&0&\cdots&0&0\\ \frac{1}{\sqrt{2}}Q&0&\frac{1}{2}I&0&\cdots&0&0\\ 0&\frac{1}{2}I&0&\frac{1}{2}I&\cdots&0&0\\ 0&0&\frac{1}{2}I&0&\ddots&\vdots&\vdots\\ \vdots&\vdots&\vdots&\ddots&\ddots&\frac{1}{2}I&0\\ 0&0&0&\cdots&\frac{1}{2}I&0&\frac{1}{\sqrt{2}}Q\\ 0&0&0&\cdots&0&\frac{1}{\sqrt{2}}Q&0\end{pmatrix}$.
Note that $G$ acts on ${\mathbb{C}}^{n}\otimes W$ as the normalized adjacency matrix of $P_{n}$, which has eigenvalues and eigenvectors
$\lambda=\cos\frac{\ell\pi}{n-1},\qquad z=\begin{pmatrix}\frac{1}{\sqrt{2}}\\ \cos\frac{\ell\pi}{n-1}\\ \cos\frac{2\ell\pi}{n-1}\\ \vdots\\ \cos\frac{(n-2)\ell\pi}{n-1}\\ \frac{(-1)^{\ell}}{\sqrt{2}}\end{pmatrix},\qquad\ell=0,1,\cdots,n-1.$
Now let $z$ be any eigenvector for $G$ with eigenvalue $\lambda$. We can decompose $z$ into $z=x+y$ where $x\in{\mathbb{C}}^{n}\otimes W$ and $y\in{\mathbb{C}}^{n}\otimes W^{\perp}$. Thus either
$\operatorname{proj}_{W}\!\left(z[\operatorname{cl}(a)]\right)=\operatorname{proj}_{W}\!\left(z[\operatorname{cl}(b)]\right)=0,$
or
$\lambda=\cos\frac{\ell\pi}{n-1},\quad\operatorname{proj}_{W}\!\left(z[\operatorname{cl}(a)]\right)=(-1)^{\ell}\operatorname{proj}_{W}\!\left(z[\operatorname{cl}(b)]\right)\neq 0.$
It follows from Theorem 8.1 that pointwise perfect $W$-transfer occurs from $a$ to $b$ at time n1n-1.
10 Discussion
We have developed some theory for perfect and pretty good subspace state transfer on simple unweighted graphs. Our characterizations connect these transport properties of walks to spectral properties of certain Hermitian weighted digraphs. An interesting direction is to explore spectral properties of the $(C_{a},C_{b})$-blow-ups, and use them to classify graph structures that admit subspace state transfer relative to specific reflection coins such as negative Grover coins and Hadamard coins. We plan on investigating these questions in the future.